OpenStudy (anonymous):

how do you do 32y^4z^3 under a sq. root sign?

OpenStudy (anonymous):

32 is 4*4*2 so sqrt 32 is 4*sqrt2 4 is 2*2 so sqrt 4 is 2 sqrt y^4=y^2 so: 2*4*y^2*sqrt(2z^3)

OpenStudy (anonymous):

sqrt(2z^3) with this part you cannot do anything

OpenStudy (anonymous):

oh but you can

OpenStudy (anonymous):

my guess is you are supposed to write this in "simplest radical form" which means you cannot have a cube under a square root.

OpenStudy (anonymous):

\[\sqrt{32y^4z^3}=\sqrt{4^2\times 2 \times y^2 \times y^2 \times z^2 \times z}\] \[=\sqrt{4^2}\times \sqrt{y^2}\times \sqrt{y^2}\times \sqrt{z^2} \times\sqrt{2z}\] \[=4y^2z\sqrt{2z}\]

OpenStudy (anonymous):

thanks, still confused a lot but thanks

OpenStudy (anonymous):

are you still confused?